# Misspecification-robust amortised simulation-based inference using variational methods

### v2

@@ 3.4 Error modelling
$\bm{\xi}(\bm{\theta})=\rm{Diag}(\rm{NN}(\bm{\theta}\,;\bm{\alpha}))+\bm{\Lambda},$ (14)
@@ 3.5 Prior over the correction model parameters
In RVNP, a (pseudo-)prior over the error model parameters $p(\bm{\xi}(\bm{\theta}))$ can be introduced to regularise the influence of the error model and make the inference more robust in the single-posterior case. In our experiments, we only use a uniform prior distribution over the error model parameters, as we focus on the robustness imparted by multiple observations.
Algorithm 1 Robust Variational Neural Posterior Estimation (RVNP)
[…]
2:repeat
3:  Sample minibatch $B\subset D$
6:until convergence
7:return $p_{\bm{\Psi}^{*}}(\bm{x}_{\rm sim}\mid\bm{\theta})$
8:
10:repeat
11:  for each $\bm{x}_{\rm obs}\in$ minibatch $\mathcal{O}$ do
13:   for $l=1$ to $K$ do
15:    $w^{(l)}\leftarrow\frac{1}{M}\sum_{m=1}^{M}\frac{p_{\bm{\xi}_{\alpha}(\bm{\theta}^{(l)})}(\bm{x}_{\rm obs}\mid\bm{x}_{\rm sim}^{(l,m)})p(\bm{\theta}^{(l)})p(\bm{\xi}_{\alpha})}{q_{\bm{\phi}}(\bm{\theta}^{(l)}\mid\bm{x}_{\rm obs})}$
16:   end for
17:   $\mathcal{L}_{V}\leftarrow-\log\left(\frac{1}{K}\sum_{l=1}^{K}w^{(l)}\right)$
19:  end for
20:until convergence
21:return $q_{\bm{\phi}^{*}}(\bm{\theta}\mid\bm{x}_{\rm obs})$, $p_{\bm{\xi}_{\alpha}^{*}(\bm{\theta})}(\bm{x}_{\rm obs}\mid\bm{x}_{\rm sim})$
22:
23:TunePosterior($q_{\bm{\phi}^{*}}$, $p_{\bm{\xi}_{\alpha}^{*}}$, $D$) // Optional: RVNP-T
24:repeat
25:  Sample $(\bm{x}_{\rm sim},\bm{\theta})\sim D$, $\bm{x}_{\rm obs}\sim p_{\bm{\xi}_{\alpha}^{*}(\bm{\theta})}(\bm{x}_{\rm obs}\mid\bm{x}_{\rm sim})$
28:until convergence
29:return $q_{\bm{\phi}^{**}}(\bm{\theta}\mid\bm{x}_{\rm obs})$
30:
31:Infer($\bm{x}_{\rm obs}$, $N_{\rm samples}$) // With optional SIR
32:for $i=1$ to $N_{\rm samples}$ do
33:  $\bm{\theta}^{(i)}\sim q_{\bm{\phi}^{*}}(\bm{\theta}\mid\bm{x}_{\rm obs})$
34:end for
36:return $\{\bm{\theta}^{(i)}\}_{i=1}^{N_{\rm samples}}$
@@ 4.1 Metrics for assessing misspecification
We consider four metrics to assess the robustness of the inference. Let $\bm{\theta}^{*}$ be the true value of the parameter. Assuming that we have a labelled test set $T=\{\bm{\theta}^{*(i)},\bm{x_{\rm obs}}^{(i)}\}_{i=1}^{N_{\rm test}}$, we evaluate the log posterior probability (LPP) of the true parameter for all points in the dataset. LPP has been extensively used in SBI literature to assess performance, and is a useful metric for assessing the ability to predict the truth from observation (2016arXiv160506376P; hermans2020likelihoodfreemcmcamortisedapproximate; ward2022robust;kelly2024misspecificationrobust; wehenkel2025addressing).
@@ 6 Discussion and limitations
In this paper, we introduced RVNP(-T), a robust amortised Bayesian inference method for simulation-based inference that jointly infers the simulation-to-reality gap and the amortised posterior using an importance-weighted autoencoder framework. RVNP is the first example of using variational inference for robust simulation-based inference, and was inspired by robust SBI approaches for synthetic likelihood (frazier2021robust), neural posterior estimation (ward2022robust), and sequential neural likelihood estimation (kelly2024misspecificationrobust). We argue that RVNP’s novelty and its ability to recover robust amortised posterior inference without setting misspecification-related tuning parameters constitutes an important step toward reliable posterior inference in amortised SBI. Previous work’s reliance on setting a hyperparameter or misspecification prior to control the influence of the misspecification on the posterior inference is undesirable, and as more data are observed we will have a better understanding of the simulation-to-reality gap, which should inform the inference without strong assumptions on the degree of misspecification. Moreover, RVNP is reliable in the well-specified case.
@@ 6.1 Error Modelling
The parameter adjustment method of kelly2024misspecificationrobust can be thought of as an error model $p_{\bm{\xi}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})=\delta(\bm{x}_{\rm obs}-\bm{x}_{\rm sim}-\bm{\xi})$ after integrating with respect to $\bm{x}_{\rm sim}$, where $\bm{\xi}$ are the linear adjustment parameters. This approach will move the observed point in summary-statistic space to a region of high probability with respect to the simulated samples. On a point-by-point basis, there are infinitely many solutions to this problem, and there is no guarantee of where an OOD point should be mapped back to in the original sample without extra information on the misspecification, such as a calibration set. Moreover, kelly2024misspecificationrobust state that the ability to recover robust posterior inference in their benchmark tasks depends on the hyperparameters that control the strength of the misspecification, which is difficult to set in practice.
@@ 6.4 Variational Inference and sample-importance-resampling
In general, we find that sample-importance-resampling is important for making the inference more robust. In particular, it is useful for "cleaning samples" that the posterior has generated which are not compatible with the prior distribution or the likelihood. We present an example of this in Figure LABEL:spectra_posterior_comparison_sir.
@@ A.5 RVNP Training Pipeline
Algorithm 2 RVNP Full Pipeline
1:Input: Flags train_flow,train_variational,tune_posteriortrain\_flow,train\_variational,tune\_posterior
2:Input: All datasets and models
3:if train_flowtrain\_flow then
4:  TrainNormalisingFlow()
5:end if
6:if train_variationaltrain\_variational then
7:  TrainVariationalPosteriorAndErrorModel()
8:end if
9:if tune_posteriortune\_posterior then
10:  PosteriorTuning()
11:end if
Algorithm 3 Train Normalising Flow Likelihood
1:Input: Simulator dataset $D=\{(\mathbf{x}_{\rm sim},\bm{\theta})\}$, normalising flow $p_{\Psi}$
2:Input: Number of epochs $E_{\rm flow}$, optimizer $\text{Optimizer}_{\Psi}$
3:for epoch =1=1 to $E_{\rm flow}$ do
4:  Sample minibatch $B_{\rm sim}\subset D$
5:  Compute NLE loss:
$\mathcal{L}_{\Psi}=-\mathbb{E}_{B_{\rm sim}}[\log p_{\Psi}(\mathbf{x}_{\rm sim}\mid\bm{\theta})]$
6:  Update $\Psi\leftarrow\text{Optimizer}_{\Psi}(\nabla_{\Psi}\mathcal{L}_{\Psi})$
7:end for
Algorithm 4 Train Variational Posterior and Error Model (RVNP)
1:Input: Observed dataset $\mathcal{O}$, normalising flow $p_{\Psi}$, posterior $q_{\bm{\phi}}$, error model network $\bm{\xi}_{\alpha}$
2:Input: Number of epochs $E_{\rm var}$, MC samples $K$, optimizers $\text{Optimizer}_{\phi},\text{Optimizer}_{\alpha}$
3:for epoch =1=1 to $E_{\rm var}$ do
4:  for each $\mathbf{x}_{\rm obs}^{(i)}\in\mathcal{O}$ do
5:   Sample $\bm{\theta}^{(l)}\sim q_{\bm{\phi}}(\bm{\theta}\mid\mathbf{x}_{\rm obs}^{(i)}),\;l=1..K$
6:   For each $\bm{\theta}^{(l)}$, sample $\mathbf{x}_{\rm sim}^{(l,m)}\sim p_{\Psi}(\mathbf{x}_{\rm sim}\mid\bm{\theta}^{(l)})$
7:   Compute IWAE variational loss:
$\mathcal{L}_{V}^{(i)}=-\log\frac{1}{K}\sum_{l=1}^{K}\frac{\mathbb{E}_{\mathbf{x}_{\rm sim}^{(l,m)}}\big[p_{\bm{\xi}_{\alpha}}(\mathbf{x}_{\rm obs}^{(i)}\mid\mathbf{x}_{\rm sim}^{(l,m)})\big]\,p(\bm{\theta}^{(l)})\,p(\bm{\xi}_{\alpha}(\bm{\theta}^{(l)}))}{q_{\bm{\phi}}(\bm{\theta}^{(l)}\mid\mathbf{x}_{\rm obs}^{(i)})}$
8:  end for
9:  Update $\phi,\alpha$ via $\text{Optimizer}_{\phi},\text{Optimizer}_{\alpha}$
10:end for
Algorithm 5 Posterior Tuning (RVNP-T)
1:Input: Simulator dataset $D$, posterior $q_{\bm{\phi}}$, error model $\bm{\xi}_{\alpha}$ (fixed)
2:Input: Number of epochs $E_{\rm tune}$, optimizer $\text{Optimizer}_{\phi}$
3:for epoch =1=1 to $E_{\rm tune}$ do
4:  Sample minibatch $B_{\rm sim}\subset D$
5:  Compute NPE-style loss with fixed error model:
$\mathcal{L}_{\rm NPE}=-\mathbb{E}_{(\mathbf{x}_{\rm sim},\bm{\theta})\in B_{\rm sim}}\mathbb{E}_{p_{\bm{\xi}_{\alpha}}(\mathbf{x}_{\rm obs}\mid\mathbf{x}_{\rm sim})}[\log q_{\bm{\phi}}(\bm{\theta}\mid\mathbf{x}_{\rm sim})]$
6:  Update $\phi\leftarrow\text{Optimizer}_{\phi}(\nabla_{\phi}\mathcal{L}_{\rm NPE})$
7:end for
@@ A.6 Training Procedure
Neural Statistic Estimator To train the neural statistic estimator on the InfoMax objective, we adopt a neural statistic and a discriminator model. We use the same optimiser parameters and batch size for the main training routine. All models are implemented in Equinox and trained using JAX. The encoder outputs a deterministic latent representation $z$ without variational sampling, and the discriminator maximises the mutual information between $z$ and $\bm{\theta}$. The spectra encoder uses one-dimensional convolutional feature extraction and global attention modelling using a Conformer block. The hidden dimension for both the embeddings and the discriminator is 100. We describe the algorithm in 6.

### v3

@@ 3.4 Error modelling
$\bm{\xi}(\bm{\theta}\,;\,\bm{\alpha}))=\rm{Diag}(\rm{NN}(\bm{\theta}\,;\bm{\alpha}))+\bm{\Lambda},$ (15)
@@ 3.5 Prior over the correction model parameters
In RVNP, a prior over the error model parameters $p(\bm{\alpha})$ can be introduced to regularise the influence of the error model parameters. In our experiments, we assume $\log p(\bm{\alpha})=0$ over the error model parameters.
@@ 4.1 Metrics for assessing misspecification
We consider four metrics to assess the robustness of the inference. Let $\bm{\theta}^{*}$ be the true value of the parameter. Assuming that we have a labelled test set $T=\{\bm{\theta}^{*(i)},\bm{x_{\rm obs}}^{(i)}\}_{i=1}^{N_{\rm test}}$, we evaluate the log posterior probability (LPP) of the true parameter for all points in the dataset. LPP has been extensively used in SBI literature to assess performance, and is a useful metric for assessing the ability to predict the truth from observation (2016arXiv160506376P; hermans2020likelihoodfreemcmcamortizedapproximate; ward2022robust;kelly2024misspecificationrobust; wehenkel2025addressing).
@@ 6 Discussion and limitations
In this paper, we introduced RVNP(-T), a robust amortised Bayesian inference method for simulation-based inference that jointly infers the simulation-to-reality gap and the amortised posterior using an importance-weighted autoencoder framework. RVNP is the first example of using variational inference for robust simulation-based inference, and was inspired by robust SBI approaches for synthetic likelihood (frazier2021robust), neural posterior estimation (ward2022robust), and sequential neural likelihood estimation (kelly2024misspecificationrobust). We argue that RVNP’s novelty and its ability to recover robust amortized posterior inference without setting misspecification-related tuning parameters constitutes an important step toward reliable posterior inference in amortised SBI. Previous work’s reliance on setting a hyperparameter or misspecification prior to control the influence of the misspecification on the posterior inference is undesirable, and as more data are observed we will have a better understanding of the simulation-to-reality gap, which should inform the inference without strong assumptions on the degree of misspecification. Moreover, RVNP is reliable in the well-specified case.
@@ 6.1 Error Modelling
The parameter adjustment method of kelly2024misspecificationrobust can be thought of as an error model $p_{\bm{\alpha}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})=\delta(\bm{x}_{\rm obs}-\bm{x}_{\rm sim}-\bm{\alpha})$ after integrating with respect to $\bm{x}_{\rm sim}$, where $\bm{\alpha}$ are the linear adjustment parameters. This approach will move the observed point in summary-statistic space to a region of high probability with respect to the simulated samples. On a point-by-point basis, there are infinitely many solutions to this problem, and there is no guarantee of where an OOD point should be mapped back to in the original sample without extra information on the misspecification, such as a calibration set. Moreover, kelly2024misspecificationrobust state that the ability to recover robust posterior inference in their benchmark tasks depends on the hyperparameters that control the strength of the misspecification, which is difficult to set in practice.
@@ 6.4 Variational Inference and sample-importance-resampling
In general, we find that sample-importance-resampling is important for making the inference more robust. In particular, it is useful for "cleaning samples" that the posterior has generated which are not compatible with the prior distribution or the likelihood. We present an example of this in Figure 7.
@@ A.5 RVNP Training Pipeline
Algorithm 1 Robust Variational Neural Posterior Estimation (RVNP)
[…]
2:repeat
3:  Sample minibatch $B\subset D$
6:until convergence
7:return $p_{\bm{\Psi}^{*}}(\bm{x}_{\rm sim}\mid\bm{\theta})$
8:
10:repeat
11:  for each $\bm{x}_{\rm obs}\in$ minibatch $\mathcal{O}$ do
13:   for $l=1$ to $K$ do
15:    $w^{(l)}\leftarrow\frac{1}{M}\sum_{m=1}^{M}\frac{p_{\bm{\alpha}}(\bm{x}_{\rm obs}\mid\bm{x}_{\rm sim}^{(l,m)},\bm{\theta}^{(l)})p(\bm{\theta}^{(l)})}{q_{\bm{\phi}}(\bm{\theta}^{(l)}\mid\bm{x}_{\rm obs})}$
16:   end for
17:   $\mathcal{L}_{V}\leftarrow-\log\left(\frac{1}{K}\sum_{l=1}^{K}w^{(l)}\right)-\log p(\bm{\alpha)}$
19:  end for
20:until convergence
21:return $q_{\bm{\phi}^{*}}(\bm{\theta}\mid\bm{x}_{\rm obs})$, $p_{\bm{\alpha}^{*}}(\bm{x}_{\rm obs}\mid\bm{x}_{\rm sim},\bm{\theta})$
22:
23:TunePosterior($q_{\bm{\phi}^{*}}$, $p_{\alpha^{*}}$, $D$) // Optional: RVNP-T
24:repeat
25:  Sample $(\bm{x}_{\rm sim},\bm{\theta})\sim D$, $\bm{x}_{\rm obs}\sim p_{\bm{\alpha}^{*}}(\bm{x}_{\rm obs}\mid\bm{x}_{\rm sim},\bm{\theta})$
28:until convergence
29:return $q_{\bm{\phi}^{**}}(\bm{\theta}\mid\bm{x}_{\rm obs})$
30:
31:Infer($\bm{x}_{\rm obs}$, $N_{\rm samples}$) // With optional SIR
32:for $i=1$ to $N_{\rm samples}$ do
33:  $\bm{\theta}^{(i)}\sim q_{\bm{\phi}^{*}}(\bm{\theta}\mid\bm{x}_{\rm obs})$
34:end for
36:return $\{\bm{\theta}^{(i)}\}_{i=1}^{N_{\rm samples}}$
@@ A.6 Training Procedure
Neural Statistic Estimator To train the neural statistic estimator on the InfoMax objective, we adopt a neural statistic and a discriminator model. We use the same optimiser parameters and batch size for the main training routine. All models are implemented in Equinox and trained using JAX. The encoder outputs a deterministic latent representation $z$ without variational sampling, and the discriminator maximises the mutual information between $z$ and $\bm{\theta}$. The spectra encoder uses one-dimensional convolutional feature extraction and global attention modelling using a Conformer block. The hidden dimension for both the embeddings and the discriminator is 100. We describe the algorithm in 2.